(Annuity) An investor needs $200,000 in 10 years. If they can get 6% interest compounded monthly, how much do they need to invest every month to achieve this total?
Answer by Cromlix(4381)

Hi there,
6% interest
100% + 6% = 106%
106% = 1.06
10 years = 120 months
(X is the unknown monthly figure)
X x 1.06^120 = 200,000
X = 200,000/1.06^120
X = 183.80
Sum to invest monthly = $184
$183.80 x 1.06^120 = $200008.91
Hope this helps :-)